OpenStudy (bloomlocke367):

@bibby. for \(A(t)=P * e^{rt}\), r is the ANNUAL interest rate, so if they give a monthly interest rate, say 1% every 3 months, would I multiply it by 4 to find the annual interest rate? (this may be a dumb question)

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

but quite frankly, r is a continuous rate of change

OpenStudy (bloomlocke367):

it says that r is interest rate (annual)

OpenStudy (amistre64):

when doing the usual compound interest calculations A = P(1+r/n)^(nt) r is the APR, n is the number of times a year its applied, and t is the number of years as n approaches infinity, we get a compounding rate that is always occuring, in infinite number of times a year as n approaches infinity, (1+r/n)^(nt) is equated as e^(rt)

OpenStudy (bloomlocke367):

the continuously compounded interest equation is \(A(t)=P * e^{rt}\), well that's what my textbook says.

OpenStudy (amistre64):

thats what i just said ... i just explained why it is that.

OpenStudy (bloomlocke367):

ohhh okay XD that just kind of confused me.

OpenStudy (bloomlocke367):

wait, I just read it like 3 more times so now I understand.

OpenStudy (amistre64):

\[\lim_{n\to\infty}~P(1+\frac rn)^{nt}=Pe^{rt}\]

OpenStudy (bloomlocke367):

Thank you.

OpenStudy (amistre64):

your welcome